### v1

# From Cradle to Cane: A Two-Pass Framework for High-Fidelity Lifespan Face Aging

Abstract
Face aging has become a crucial task in computer vision, with applications ranging from entertainment to healthcare. However, existing methods struggle with achieving a realistic and seamless transformation across the entire lifespan, especially when handling large age gaps or extreme head poses. The core challenge lies in balancing age accuracy and identity preservation—what we refer to as the Age-ID trade-off. Most prior methods either prioritize age transformation at the expense of identity consistency or vice versa. In this work, we address this issue by proposing a two-pass face aging framework, named Cradle2Cane, based on few-step text-to-image (T2I) diffusion models. The first pass focuses on solving age accuracy by introducing an adaptive noise injection (AdaNI) mechanism. This mechanism is guided by including prompt descriptions of age and gender for the given person as the textual condition. Also, by adjusting the noise level, we can control the strength of aging while allowing more flexibility in transforming the face. However, identity preservation is weakly ensured here to facilitate stronger age transformations. In the second pass, we enhance identity preservation while maintaining age-specific features by conditioning the model on two identity-aware embeddings (IDEmb): SVR-ArcFace and Rotate-CLIP. This pass allows for denoising the transformed image from the first pass, ensuring stronger identity preservation without compromising the aging accuracy. Both passes are jointly trained in an end-to-end way. Extensive experiments on the CelebA-HQ test dataset, evaluated through Face++ and Qwen-VL protocols, show that our Cradle2Cane outperforms existing face aging methods in age accuracy and identity consistency. Additionally, Cradle2Cane demonstrates superior robustness when applied to in-the-wild human face images, where prior methods often fail. This significantly broadens its applicability to more diverse and unconstrained real-world scenarios.
1 Introduction
Deep learning lecun2015deeplearning has allowed a realistic alteration of the apparent age of a person chen2023face ; deng2019arcface ; zhang2017age , opening promising applications in areas such as computer graphics, entertainment, forensics and healthcare. The goal of facial age transformation is to simulate the natural aging or de-aging process in a visually convincing manner. To this end, numerous methods have been developed to achieve high-quality, identity-preserving age progression and regression. Recent approaches are based on deep generative models, such as generative adversarial networks (GANs) alaluf2021only ; goodfellow2014gan ; hsu2022agetransgan and Diffusion Models (DMs) banerjee2023identity ; chen2023face ; ito2025selfage ; wahid2024diffage3d , and have shown promising results. However, to the best of our knowledge, most existing methods suffer from a limited transformation range and often struggle to maintain high-fidelity results when handling large age gaps, occlusions, or extreme head poses. As a result, they fall short of delivering seamless cradle-to-cane face aging performance across the entire lifespan.
In this study, we attribute the limitations of existing face aging methods to an imbalanced trade-off between age accuracy and identity consistency—a challenge we term the Age-ID trade-off. Most prior approaches alaluf2021only ; chen2023face ; ito2025selfage tend to emphasize one aspect while neglecting the other due to their unified framework to deal with entire lifespan ages, resulting in either visually convincing age transformations that compromise identity, or identity-preserving outputs with inaccurate aging effects. This imbalance is evident in the trade-off curves shown in Fig. 1, where, for example, as the age difference increases, methods such as FADING chen2023face , CUSP gomez2022custom , and Lifespan or2020lifespan tend to show improved aging realism at the cost of reduced identity preservation, or vice versa. The fluctuating curves of harmonic means further demonstrate this phenomenon. To address the Age-ID trade-off problem, we propose our face aging framework built upon few-step T2I diffusion models SDXL-Turbo sauer2024adversarial , which offer two key advantages: 1) the few-step nature of these models dao2024swiftbrushv2 ; li2023faster ; liu2023instaflow ; sauer2024adversarial enables fast inference while maintaining high image fidelity, and 2) the flexible noise control in the forward diffusion process allows fine-grained modulation of aging strength by adjusting the injected noise scale. As illustrated in Fig. 2, injecting higher noise levels during the forward diffusion process increases editability, enabling more pronounced aging transformations while downgrading the identity consistency. In contrast, lower noise levels better preserve identity information with less aging accuracy, highlighting the trade-off between visual age change and identity consistency. Similar identity-editing trade-offs are also observed in image editing and generation hertz2022prompt ; li2023stylediffusion ; meng2021sdedit ; kai2023DPL . However, directly applying the few-step diffusion models cannot achieve fine-grained face aging with identity consistency and age accuracy, resulting from that the few-step T2I diffusion models ho2022imagen ; Rombach_2022_CVPR_stablediffusion ; deepfloyd do not inherently support age or identity conditions.
Figure 1: Age–ID trade-off curves across sixty age shift values. We compute the Age/ID cosine similarities over 100 human faces across 1-60 age shift values and the corresponding harmonic means. Existing approaches tend to favor either age accuracy or identity consistency, resulting in imbalanced performance across the entire lifespan ages. In contrast, our method Cradle2Cane achieves a better balance between the two objectives. More details and results are provided in Appendix C.5.
In this paper, we propose to address the Age-ID trade-off by decoupling age accuracy and identity preservation into a two-pass ***We define a pass as the T2I inference process that generates a real image through the diffusion model. diffusion framework Cradle2Cane, with SDXL-Turbo as the backbone and each stage is tailored to optimize a specific objective. During the first pass, which focuses on precise age control, we introduce an adaptive noise injection (AdaNI) mechanism guided by textual descriptions which containing age and gender attributes. The level of noise injected is dynamically adjusted based on the magnitude of the desired age transformation. Naturally, human identity is better preserved with smaller age variations and tends to degrade with larger age gaps. This strategy aims to overcome the limitations of existing methods that rely on uniform solutions for modeling aging across the entire lifespan. In this stage, identity is only weakly preserved to allow greater flexibility in age manipulation. In the second pass, we reinforce the identity consistency of the generated face while preserving the age-specific characteristics from the first pass. We propose to achieve this by conditioning the few-step diffusion model with a concatenation of two identity-aware embeddings (IDEmb): an SVR-ArcFace embedding and a Rotate-CLIP embedding. These embeddings guide the model to denoise a minimally perturbed input, ensuring stronger identity preservation without compromising the age transformation. It is worth noting that both stages are jointly trained in an end-to-end manner, where the output image from the first stage is further diffused and used as the noisy latent input for the second stage. After training, our proposed method, Cradle2Cane, is capable of achieving high-fidelity and adaptive face aging while maintaining a superior balance between age accuracy and identity consistency compared to existing approaches.
In our experiments, we conduct comprehensive comparisons against a diverse set of GAN-based and diffusion-based face aging methods on the CelebA-HQ karras2017progressive ; liu2015celeba test dataset. We adopt both the Face++ gomez2022custom and Qwen-VL qwenvl2024 evaluation protocols to assess performance in terms of age accuracy and identity consistency. Both evaluation pipelines consistently validate the effectiveness of our method, Cradle2Cane, which achieves a superior balance in the age-ID trade-off with inference speeds comparable to GAN-based methods. Furthermore, benefiting from the strong generative capacity of text-to-image diffusion models, Cradle2Cane exhibits enhanced robustness on in-the-wild human face images—where previous approaches often struggle—thereby significantly broadening its range of practical application scenarios. To summarize, this paper makes the following main contributions:
  • We propose a novel two-pass approach Cradle2Cane that decouples age accuracy and identity preservation in face aging, where the first pass applies adaptive noise injection (AdaNI) for precise age manipulation, and the second pass reinforces identity consistency through identity-aware embedding (IDEmb).
  • For the first pass, we introduce a text-guided adaptive noise injection (AdaNI) strategy that dynamically adjusts the injected noise level based on the desired age transformation strength, enabling fine-grained control over the age-ID trade-off.
  • To enhance identity preservation, we design a conditioning mechanism that leverages a combination of SVR-ArcFace and Rotate-CLIP as identity-aware embeddings (IDEmb), guiding the second-pass denoising process for high-fidelity and identity-consistent outputs.
  • Extensive experiments on the CelebA-HQ test dataset demonstrate that Cradle2Cane consistently outperforms existing baselines across age accuracy, identity consistency and image quality, while maintaining fast inference speed. Moreover, Cradle2Cane exhibits strong generalization to in-the-wild human face images—a challenging scenario where current methods often fail.
Figure 2: (Left) We illustrate the effects of injecting three different levels of noise into the input image, as used in the 4-step SDXL-Turbo image-to-image pipeline. As visually evident, higher noise levels lead to more pronounced age transformations at the cost of reduced identity preservation. (Right) We present a statistical analysis on 100 human faces, that quantitatively demonstrates the Age-ID trade-off inherent in face aging tasks. Specifically, we evaluate three representative noise injection levels and measure their corresponding impacts on age accuracy and identity consistency.
2 Related Work
Face Aging. Facial age editing aims to simulate the natural process of fine-grained aging in facial images while faithfully preserving the subject’s identity. Traditional approaches relied on physical modeling ramanathan2006modeling ; suo2012concatenational or attribute manipulation kemelmacher2014illumination ; shu2015personalized , but often struggled with generalization and photorealism. The emergence of GAN-based methods such as Lifespan or2020lifespan , IPCGAN wang2018face , and CAAE zhang2017age significantly improved aging realism by learning conditional generative mappings from large-scale datasets. For instance, SAM alaluf2021only combines an aging encoder with an inversion encoder to perform age transformations in the latent space of StyleGAN2. CUSP gomez2022custom disentangles style and content using dual encoders and manipulates them for personalized age transformations. HRFAE yao2021high introduces an age modulation network that fuses age labels into latent representations to guide high-resolution age progression. With the recent advancements in diffusion models ho2020denoising ; song2020denoising , they have emerged as powerful alternatives for high-fidelity face aging chen2023face ; ito2025selfage ; kim2022diffusionclip . For example, FADING chen2023face fine-tunes a pretrained LDM Rombach_2022_CVPR_stablediffusion on age-labeled datasets. During inference, it uses NTI hertz2022prompt ; mokady2022null to embed input images into latent space, allowing for localized age edits. Similarly, IPFE banerjee2023identity combines latent diffusion with biometric and contrastive losses to enforce identity preservation during facial aging and de-aging. However, we posit that face aging should adhere to a natural principle: subtle age variations preserve facial identity more effectively, whereas significant age gaps introduce greater identity distortion, while the current approaches generally overlook this consideration.
Semantic Latent Spaces in Generative Models. Linear latent space models of facial shape and appearance were extensively studied in the 1990s, primarily through PCA-based representations blanz2023morphable_3d ; cootes2001active_appearance_models ; rowland1995manipulating_facial . However, these early approaches were limited to aligned and cropped frontal facial images. Afterwards, the StyleGAN-family generative models karras2019stylegan1 ; karras2020analyzing_stylegan2 , have demonstrated powerful editing capabilities, largely attributed to the structured and interpretable nature of their latent spaces. In contrast, diffusion models lack an explicit latent space by design. Nevertheless, recent studies have attempted to uncover GAN-like latent structures within them, targeting various representations such as the UNet bottleneck haas2024discovering ; kwon2022diffusion_asryp ; park2023understanding ; wu2023latent , the noise input space dalva2024noiseclr , and the text embedding space baumann2024continuous_attr_ctrl . For example, Concept Sliders gandikota2024concept_slider propose semantic image editing through low-rank adaptation in weight space, guided by contrastive image or text pairs. Despite these advances, existing disentanglement-based methods are typically limited to coarse-grained attribute control—such as adjusting age, hair, or expression via semantical direction controls—and often struggle to achieve precise, fine-grained manipulation of facial aging features.
Text-to-Image Models Distillation. Text-to-image (T2I) models based on diffusion balaji2022ediffi ; dai2023emu ; rombach2022high ; saharia2022photorealistic have achieved impressive progress in generating high-quality images from text prompts. Despite their success, the inference phase remains a bottleneck—diffusion models require iterative denoising. To mitigate this, a variety of acceleration methods have been proposed. While training-free approaches have shown promise for both diffusion huang2024harmonica ; li2023faster ; lou2024token_caching ; zheng2023_fastsampling , the most effective strategies often involve additional distillation process to accelerate the sampling process beyond the capabilities of the original base models. SD-Turbo sauer2024adversarial introduces a discriminator combined with a score distillation loss to improve performance. Most of these methods depend on image-text pair datasets for training, requiring substantial data alignment between visual and textual features. In contrast, SwiftBrush nguyen2023swiftbrush adapts variational score distillation. SwiftBrush nguyen2023swiftbrush achieves the first image-free training by using generated images as the training set, avoiding the need for paired datasets. In this paper, we build our method, Cradle2Cane, upon SDXL-Turbo sauer2024adversarial , which is widely adopted and demonstrate strong performance in few-step high-quality image generation, to introduce the two-pass architecture Cradle2Cane tailored for controllable facial age transformation.
3 Method
In this section, we present our framework Cradle2Cane for face aging. Given an source face image $\mathbf{x}_{a}$ of a person at source age $a$, and a target age $b$, our goal is to generate a realistic target face image $\mathbf{x}_{b}$, depicting the same identity at age $b$. The main challenge lies in achieving realistic aging effects while preserving the identity (ID) of the subject. Due to the scarcity of datasets containing the same identity across a wide age range, directly transforming $\mathbf{x}_{a}$ to $\mathbf{x}_{b}$ remains a difficult task. The full pipeline of our method Cradle2Cane is visualized in Fig. 3. We first introduce the preliminaries in Section 3.1. Then Section 3.2 presents the adaptive noise injection (AdaNI) during the first pass. Section 3.3 details the second pass with identity-aware embedding (IDEmb) for identity preservation. Section 3.4 defines the training objectives and loss functions.
3.1 Preliminary
Fast Sampling of T2I diffusion models. SDXL-Turbo sauer2024adversarial accelerates standard diffusion models podell2023sdxl ; rombach2021highresolution via Adversarial Diffusion Distillation, enabling high-quality image generation in only a few steps. Unlike DDPM ho2020denoising or DDIM song2020denoising , which typically require 50 to 1000 inference steps, SDXL-Turbo achieves 1-4 steps generation by training a compact denoiser to imitate a large teacher model, supervised jointly by distillation and adversarial losses. The forward process perturbs an initial latent variable $\mathbf{z}_{0}\in\mathbb{R}^{d}$ into increasingly noisy states $\mathbf{z}_{1},\dots,\mathbf{z}_{T}$ using a Markov chain:
$q(\mathbf{z}_{1:T}\mid\mathbf{z}_{0})=\prod_{t=1}^{T}q(\mathbf{z}_{t}\mid\mathbf{z}_{t-1}),$ (1)
where $T$ is the denoise steps. The reverse process then reconstructs $\mathbf{z}_{0}$ from $\mathbf{z}_{T}$ in $T$ learned steps:
$p_{\theta}(\mathbf{z}_{0:T})=p(\mathbf{z}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}),$ (2)
where each $p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t})$ is a Gaussian parameterized by a neural network trained to approximate the inverse of the forward noising process.
Overall pipeline of our method Cradle2Cane. To address the challenge of controllable and identity-preserving face aging, we propose a two-pass framework, Cradle2Cane, built upon the efficient SDXL-Turbo model. In the first stage, we perform adaptive noise injection (AdaNI) on the input face image $\mathbf{x}_{a}$, guided by age-specific embedding, to generate an intermediate image $\hat{\mathbf{x}}_{b}$ that reflects the target age $b$. This step aims to synthesize realistic aging effects while maintaining essential identity traits. However, for large age gaps, $\hat{\mathbf{x}}_{b}$ may exhibit partial identity drift due to the strong age transformation. To compensate for this, the second stage focuses on enhancing identity consistency. A lower magnitude of noise is injected into $\hat{\mathbf{x}}_{b}$, and identity-aware embeddings (IDEmb) conditioning is applied using features extracted from the original input image. This results in the final output face image $\mathbf{x}_{b}$, which exhibits both faithful aging effects and high identity preservation.
3.2 1st Pass: Adaptive Noise Injection (AdaNI) for Age Accuracy
We address the Age-ID trade-off by focusing on two critical aspects: age accuracy and identity preservation. Empirically, we observe that the extent of facial modifications required during age progression is closely correlated with the magnitude of the age gap. Specifically, larger age transitions typically demand more pronounced structural and textural changes, while smaller transitions involve only minor appearance adjustments. Prior works meng2021sdedit , as well as the left portion of Fig. 2, suggest that the level of noise injected into the input image controls the flexibility and intensity of editing.
Building on this insight, we conduct a systematic study to examine how varying noise levels influence the balance between identity fidelity and age realism. In particular, we apply three levels of noise injection, denoted as $z_{1}$, $z_{2}$, and $z_{3}$, to a set of 100 face images under age transformation tasks spanning age shifts from 1 to 60 years. As illustrated in Fig. 2 (right), lower noise injection intensity ($z_{1}$) consistently leads to superior identity preservation across all age shifts. However, it fails to deliver accurate age progression, particularly for larger age gaps. Conversely, higher noise levels ($z_{3}$) produce more realistic age transformations but significantly compromise identity consistency. These results highlight a clear trade-off between identity preservation and age accuracy, governed by the noise injection intensity. Motivated by these findings, we propose an adaptive noise injection (AdaNI) strategy that dynamically modulates the noise level based on the target age shift.
More specifically, we encode a predefined age prompt using the CLIP text encoder to obtain the text embedding, which conditions the generation process via cross-attention. For AdaNI injection, we divide the age transformation magnitude into three categories, using ages 5 and 20 as boundaries based on our quantitative analysis in Fig. 2, where age accuracy drops significantly beyond these thresholds. Each category corresponds to a specific noise level applied during first-pass noise injection:
$\hat{\mathbf{z}}_{0}=\begin{cases}p_{\theta}(\mathbf{z}_{0}\mid\mathbf{z}_{1}),&|\Delta\text{age}|\leq 5,\\ p_{\theta}(\mathbf{z}_{0}\mid\mathbf{z}_{2}),&5<|\Delta\text{age}|\leq 20,\\ p_{\theta}(\mathbf{z}_{0}\mid\mathbf{z}_{3}),&|\Delta\text{age}|>20,\end{cases}$ (3)
where $\mathbf{z}_{1},\mathbf{z}_{2},\mathbf{z}_{3}$ represent different noise levels injected into the latent space. After completing the diffusion process to obtain the final latent code $\hat{\mathbf{z}}_{0}$, the intermediate aged face is reconstructed via the VAE decoder $D$: $\hat{\mathbf{x}}_{b}=D(\hat{\mathbf{z}}_{0})$, which exhibits high age accuracy but relatively weak identity preservation. This enables our model to better balance the competing objectives of age accuracy and faithful identity preservation. Nonetheless, even with adaptive injection, identity degradation becomes increasingly prominent with larger age transitions. To mitigate this effect, the second pass of Cradle2Cane explicitly enhances identity consistency by refining identity-specific embeddings.
Figure 3: Our method Cradle2Cane consists of two passes: the first pass employs adaptive noise injection (AdaNI) to enhance age accuracy, while the second pass incorporates identity-aware embeddings (IDEmb), including SVR-ArcFace and Rotate-CLIP embeddings, to improve identity consistency. During training the MLPs and UNet-LoRA modules, we jointly optimize identity loss between source and target face images, as well as age and quality losses over the target images.
3.3 2nd Pass: Identity-Aware Embedding (IDEmb) for Identity Preservation
To further improve identity preservation, we extract identity-aware embeddings (IDEmb) from the source face $\mathbf{x}_{a}$ using both ArcFace and CLIP encoders, which are standard features alaluf2021only ; banerjee2023identity ; shen2023clip for measuring and guiding identity information. A central challenge in this process is the inherent entanglement between age and identity within these embeddings—both ArcFace and CLIP features tend to encode age-related cues alongside identity information guerrero2024texsliders ; shiohara2023blendface . To overcome this limitation, we propose two novel embedding modules: SVR-ArcFace and Rotate-CLIP. These modules are designed to explicitly suppress age-related components within their respective embedding spaces, thereby disentangling identity from age.
3.3.1 SVR-ArcFace
Given a source face image $\mathbf{x}_{a}$, we generate a set of $n$ aged face images $\{\mathbf{x}_{b}^{(i)}\}_{i=1}^{n}$ by injecting different noise levels in the first stage. These images share the same identity as $\mathbf{x}_{a}$ but exhibit different age characteristics. Inspired by prior works gu2014weighted ; li2023get ; liu2025onepromptonestory , which suggest that applying Singular Value Decomposition (SVD) followed by singular value reweighting (SVR) can enhance shared features while suppressing divergent ones such as age, we propose a singular value reweighting technique to refine identity features from the ArcFace embeddings. We refer to this method as SVR-ArcFace.
First, we extract ArcFace embeddings $u_{a}$ and $\{u_{b}^{(i)}\}_{i=1}^{n}$ from the source and aged face images, and concatenate them into a matrix:
$U=[u_{a},u_{b}^{(1)},u_{b}^{(2)},\ldots,u_{b}^{(n)}]\in\mathbb{R}^{D\times(n+1)},$ (4)
where $D$ is the embedding dimension. We then perform Singular Value Decomposition (SVD) on $U$:
$U=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{T},\quad\boldsymbol{\Sigma}=\mathrm{diag}(\sigma_{0},\sigma_{1},\ldots,\sigma_{r}),$ (5)
where $r=\min(D,n+1)$. Following the assumption in previous works, we treat the dominant singular values of $U$ as encoding the shared identity, since all embeddings in $U$ correspond to the same person. To suppress age-related variations and emphasize identity features, we apply a nonlinear function for the singular value reweighting (SVR):
$\hat{\sigma}_{i}=\beta e^{\alpha\sigma_{i}}\cdot\sigma_{i},$ (6)
where $\alpha,\beta>0$ are hyperparameters that control the enhancement strength. The reweighted singular value matrix is defined as $\hat{\boldsymbol{\Sigma}}=\mathrm{diag}(\hat{\sigma}_{0},\hat{\sigma}_{1},\ldots,\hat{\sigma}_{r})$, and then the refined embedding matrix is reconstructed as:
$\hat{U}=\mathbf{U}\hat{\boldsymbol{\Sigma}}\mathbf{V}^{T}.$ (7)
Finally, we use the first column of $\hat{U}$, denoted as $\hat{u}_{a}$, as the refined identity embedding to guide the identity preservation in the second stage.
3.3.2 Rotate-CLIP
Given a source face image $\mathbf{x}_{a}$ with source age $a$ and target age $b$, we extract the CLIP image embedding $i_{a}=I_{\text{CLIP}}(\mathbf{x}_{a})$, along with the text embeddings $t_{a}=T_{\text{CLIP}}(a)$ and $t_{b}=T_{\text{CLIP}}(b)$, using the pretrained CLIP image encoder $I_{\text{CLIP}}(\cdot)$ and text encoder $T_{\text{CLIP}}(\cdot)$. Our goal is to shift the age-related component in $i_{a}$ toward the target age domain in CLIP space, leveraging CLIP’s joint visual-textual alignment. A common approach, inspired by shen2023clip , is to compute the age shift vector as the difference of text embeddings:
$\Delta=t_{b}-t_{a}.$ (8)
However, this simple subtraction may introduce semantic inconsistencies due to CLIP’s coarse age representations wolfe2022markedness ; zarei2024understanding . To address this, we propose a rotational projection using spherical linear interpolation (slerp), which more smoothly captures semantic transitions between ages:
$\Delta^{\prime}=\operatorname{slerp}(t_{b},t_{a},\lambda),$ (9)
where $\lambda\in[0,1]$ controls the interpolation. The Rotate-CLIP embedding is then defined as:
$\hat{i}_{a}=i_{a}+\Delta^{\prime},$ (10)
which shifts $i_{a}$ toward the target age direction while preserving other identity-related information.
The refined identity embeddings $\hat{u}_{a}$ and $\hat{i}_{a}$, obtained from SVR-ArcFace and Rotate-CLIP, are projected through two MLPs to align with the text-embedding feature dimension, then concatenated to form IDEmb before injected into the cross-attention module of SDXL-Turbo:
$\tilde{u}_{a}=\operatorname{MLP}_{u}(\hat{u}_{a}),\quad\tilde{i}_{a}=\operatorname{MLP}_{i}(\hat{i}_{a}).$ (11)
3.4 Training Losses
Based on the architecture described above, we jointly optimize the MLPs and UNet-LoRA hu2022lora ; simoLoRA2023 modules using a weighted combination of three objectives: identity loss, age loss, and quality loss. The ArcFace and CLIP encoders remain frozen during training.
Identity Loss.
To preserve facial identity, we employ a combination of multi-scale structural similarity (MS-SSIM) wang2003multiscale and high-level identity embedding similarity. Specifically, we use a pretrained ArcFace deng2019arcface model to extract embeddings and compute the cosine distance between the source image $\mathbf{x}_{a}$ and the generated image $\mathbf{x}_{b}$:
$\mathcal{L}_{\text{id}}=\lambda_{1}\cdot\left(1-\mathrm{MS\text{-}SSIM}(\mathbf{x}_{a},\mathbf{x}_{b})\right)+\lambda_{2}\cdot\left(1-\cos\left(f_{\mathrm{Arc}}(\mathbf{x}_{a}),f_{\mathrm{Arc}}(\mathbf{x}_{b})\right)\right),$ (12)
where $f_{\mathrm{Arc}}(\cdot)$ denotes the ArcFace encoder.
Age Loss.
To ensure age accuracy, we define an age loss that measures both visual consistency and numerical correctness. The first term computes the cosine similarity between embeddings of the intermediate result $\hat{\mathbf{x}}_{b}$ and the final output $\mathbf{x}_{b}$ using a pretrained MiVOLO kuprashevich2023mivolo model. The second term minimizes the L2 distance between the predicted and target ages:
$\mathcal{L}_{\text{age}}=\lambda_{3}\cdot\left(1-\cos\left(f_{\mathrm{Mi}}(\hat{\mathbf{x}}_{b}),f_{\mathrm{Mi}}(\mathbf{x}_{b})\right)\right)+\lambda_{4}\cdot\left\|g_{\mathrm{Mi}}(\mathbf{x}_{b})-b\right\|_{2}^{2},$ (13)
where $f_{\mathrm{Mi}}(\cdot)$ and $g_{\mathrm{Mi}}(\cdot)$ denote the MiVOLO feature extractor and age estimator, respectively.
Quality Loss.
To improve perceptual fidelity, we combine the LPIPS metric zhang2018unreasonable , which measures perceptual similarity aligned with human vision, with an adversarial loss from a GAN goodfellow2014gan discriminator to encourage photorealism:
$\mathcal{L}_{\text{per}}=\lambda_{5}\cdot\mathrm{LPIPS}(\mathbf{x}_{a},\mathbf{x}_{b})+\lambda_{6}\cdot\mathcal{L}_{\mathrm{GAN}}(\mathbf{x}_{b}).$ (14)
Overall Objective.
The final training objective is a weighted sum of the three losses:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{id}}+\mathcal{L}_{\text{age}}+\mathcal{L}_{\text{per}},$ (15)
where $\lambda_{1}$ through $\lambda_{6}$ are scalar coefficients that balance the contributions of each component.
4 Experiments
4.1 Experimental Setups
Evaluation Benchmarks and Metrics. We evaluate our method on two datasets: a subset of CelebA-HQ karras2017progressive and CelebA-HQ (in-the-wild) test datasets. For each dataset, we randomly select 100 face images per gender. Each image is used to generate age-progressed faces from 0 to 80 years in 5-year intervals, resulting in 3,200 test images per dataset. We also use Carvekit CarveKit to remove background. Following prior works gomez2022custom ; yao2021high , we utilize the Face++ API to quantitatively assess age estimation accuracy, identity preservation, and image quality. In addition, we employ large multimodal models, such as Qwen-VL qwenvl2024 , to conduct high-level perceptual evaluations. These models provide interpretable assessments of perceived age, identity consistency, and visual realism via carefully designed task-specific prompts. To jointly evaluate age accuracy and identity preservation, we propose the Harmonic Consistency Score (HCS), a unified metric that balances both factors. Full metric definitions and evaluation prompt templates are provided in Appendix C.
Comparison Methods. To evaluate the performance of our method, we compare it with several state-of-the-art face aging baselines. Specifically, we include: (1) Diffusion-based methods: IPFE banerjee2023identity , FADING chen2023face ; and (2) GAN-based methods: SAM alaluf2021only , CUSP gomez2022custom , Lifespan or2020lifespan , and HRFAE yao2021high . Detailed configurations and implementations of both our method and baselines are included in Appendix C.
Figure 4: Qualitative comparison with existing face aging methods across lifespan ages. Our method Cradle2Cane is even able to imitate the natural hair change while the previous methods cannot. For comparisons on in-the-wild images, please refer to Fig. 8 in the Appendix.
Figure 5: (Left) While applying to in-the-wild real human faces, Cradle2Cane demonstrates better performance while the existing methods often fail. (Right) Our Cradle2Cane can also be applied to modify gender and emotion attributes while performing age transformation on human faces.
4.2 Experimental Results
Quantitative Comparison. As demonstrated in Table 2, Cradle2Cane consistently outperforms existing face aging methods in both the Face++ and Qwen-VL evaluations across the CelebA-HQ and CelebA-HQ (in-the-wild) datasets. It achieves the lowest age estimation error, the highest image quality scores, and the best HCS values, while maintaining competitive identity preservation. Notably, Cradle2Cane achieves these results with a relatively small training set (10K) and a fast inference time (0.56s). These results underscore the effectiveness and efficiency of our framework in balancing aging realism, identity consistency, and visual quality across diverse evaluation protocols.
Qualitative Comparison. Figure 4 presents a visual comparison of face aging results between Cradle2Cane and recent GAN- and diffusion-based baselines. Compared to other methods, our approach demonstrates more realistic aging transitions with consistent identity preservation across all age ranges. In contrast, prior methods often exhibit texture artifacts, age realism issues, or identity shifts, particularly at extreme ages. Our method, however, produces natural skin aging, hair graying, and structural changes, reflecting a superior modeling of facial aging patterns. These results emphasize the visual fidelity and robustness of our framework.
Ablation Study. We conduct an ablation study to assess the impact of each proposed component, as shown in Table 2. Removing our aging mechanism AdaNI results in a substantial age estimation error (±8.87), emphasizing its critical role in achieving age accuracy. Introducing AdaNI alone significantly reduces the error (±3.94), though it slightly compromises identity similarity. Incorporating the SVR-ArcFace module improves identity consistency (from 59.70 to 63.38), validating its effectiveness for identity preservation. Finally, adding Rotate-Clip further enhances identity performance and contributes to a well-balanced trade-off across all metrics. Notably, the overall HCS score steadily increases throughout, with the full configuration achieving the highest score (78.33).
Table 1: Quantitative comparison using both Face++ and Qwen-VL evaluation protocols on CelebA-HQ and CelebA-HQ (in-the-wild) test dataset. We calculate the age accuracy, identity preservation, image quality and the Harmonic consistency score (HCS) to compare with existing face aging methods. Best results are marked in blue, and second-best in green.
Method Type Face++ Evaluation (CelebA-HQ) Qwen-VL Evaluation (CelebA-HQ) Qwen-VL (CelebA-HQ-in-the-wild) Inference Time (s) Train Data
Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑ Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑ Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑
Lifespan or2020lifespan GAN ±22.07 79.80 66.68 57.40 ±27.99 71.99 86.03 42.38 ±26.20 71.14 69.41 46.47 0.95 70K
HRFAE yao2021high GAN ±15.12 94.32 62.28 74.95 ±17.77 77.86 90.93 62.19 ±19.98 77.68 84.49 60.87 0.17 300K
SAM alaluf2021only GAN ±8.42 81.96 68.38 80.42 ±6.31 72.15 90.70 77.72 ±6.86 54.87 87.10 66.01 0.39 70K
CUSP gomez2022custom GAN ±9.59 85.92 64.98 80.67 ±7.45 74.44 88.06 77.84 ±13.66 76.89 81.86 70.94 0.24 30K
FADING chen2023face Diffusion ±14.47 86.70 64.65 73.52 ±7.90 75.08 90.02 77.57 ±9.25 73.33 88.01 75.06 61.26 -
IPFE banerjee2023identity Diffusion ±11.95 75.14 63.55 72.54 ±11.67 69.40 87.01 70.01 ±12.97 65.34 88.03 66.43 8.84 -
Cradle2Cane Diffusion ±7.47 81.34 72.69 81.33 ±4.62 70.29 92.37 78.33 ±5.05 67.15 88.92 75.94 0.56 10K
Table 2: Ablating each component with Qwen-VL evaluation.
AdaNI SVR-ArcFace Rotate-Clip Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑
$\times$ $\times$ $\times$ ±8.87 68.92 92.00 73.10
$\checkmark$ $\times$ $\times$ ±3.94 59.70 92.15 71.83
$\times$ $\checkmark$ $\times$ ±9.48 70.17 92.16 73.11
$\checkmark$ $\checkmark$ $\times$ ±6.75 63.38 92.43 71.92
$\checkmark$ $\checkmark$ $\checkmark$ ±4.62 70.29 92.37 78.33
Figure 6: User Study
Ours vs CUSPOurs vs SAM0025252525505050507575757510010010010072727272676767672828282833333333
User Study. To assess human-perceived quality, we conduct a user study comparing our method Cradle2Cane with two state-of-the-art face aging methods: SAM alaluf2021only and CUSP gomez2022custom . We randomly sample 20 identity images from the CelebA-HQ test set and generate 6 aging results for each, evenly spaced from age 5 to 80. 50 volunteers are asked to perform pairwise comparisons between our results and each baseline, considering three joint criteria—age accuracy, identity preservation, and overall image quality. Each query follows a forced 1-vs-1 protocol with randomized display order to prevent position bias. As summarized in Fig. 6, our method is consistently preferred by a clear majority, demonstrating superior perceptual quality and better alignment with human judgment.
Additional Applications. Since our method Cradle2Cane is build upon the large T2I diffusion model, it is also able to deal with various in-the-wild images while the previous methods fail (Fig. 5-(Left). Besides facial age transformation, our method can be easily adapted to other facial editing tasks, such as gender transformation and expression modification. As illustrated in Fig. 5-(Right), our approach achieves gender and expression changes while maintaining high identity consistency. This further demonstrates the versatility and generalizability of our facial editing framework.
5 Conclusion
In this work, we tackle the fundamental challenge of achieving both age accuracy and robust identity preservation in face aging—a problem we term the Age-ID trade-off. While existing methods often prioritize one objective at the expense of the other, our proposed framework, Cradle2Cane, introduces a two-pass framework that explicitly decouples these goals. By leveraging the flexibility of few-step text-to-image diffusion models, we introduce an adaptive noise injection (AdaNI) mechanism for fine-grained age control in the first pass, and reinforce identity consistency through dual identity-aware embeddings (IDEmb) in the second pass. Our method is trained end-to-end, enabling high-fidelity, controllable age transformation across the full lifespan, while significantly improving inference speed and visual realism. Extensive evaluations on CelebA-HQ confirm that Cradle2Cane achieves new state-of-the-art performance in terms of both age accuracy and identity preservation. In addition, Cradle2Cane demonstrates strong generalization to real-world scenarios by effectively handling in-the-wild human face images, a setting where existing methods often fail.
Appendix
Appendix A Limitations
While our method achieves state-of-the-art performance in balancing age realism and identity consistency, there remain several limitations that merit discussion. In cases of extreme age transformation (e.g., from a child to an elderly person or vice versa), the model tends to favor facial realism and age accuracy at the cost of preserving some visual details in the original image. For example, accessories such as eyeglasses, earrings, or clothing color may not always be faithfully retained after editing, as they are not explicitly modeled or enforced during training. This issue stems from the adaptive noise injection design, which purposefully increases editability for large age gaps, potentially altering finer image semantics beyond facial identity.
Appendix B Broader Impacts
Our proposed face aging method provides a flexible framework for independently controlling visual age via a two-pass diffusion process. This enables a range of positive applications, including digital entertainment (e.g., age effects in movies or games), age-invariant face recognition, and future appearance simulation for healthcare or counseling. All experiments are conducted on anonymized public benchmarks under ethical research settings.
At the same time, we recognize the potential risks associated with misuse. The ability to generate photorealistic age manipulation with identity consistency may facilitate malicious uses such as identity spoofing, misinformation, or privacy violations. We strongly discourage unauthorized or commercial deployment without safeguards like watermarking, traceable provenance, or human review. We hope this work inspires further progress toward ethical and responsible generative modeling in the vision community.
Appendix C Experiments details
C.1 Implementation Details.
[…]
The U-Net backbone of SDXL-Turbo is fine-tuned via LoRA, and training is conducted on 8 NVIDIA A6000 GPUs with a batch size of 4. For hyperparameters, in SVR-ArcFace, we set $\alpha=0.01$ and $\beta=1.2$, while in Rotate-CLIP, we use an interpolation strength of $\lambda=0.5$. For the training loss terms, we set the weights as $\lambda_{1}=0.25$, $\lambda_{2}=1.2$, $\lambda_{3}=1.5$, $\lambda_{4}=1.5$, $\lambda_{5}=0.25$, and $\lambda_{6}=0.1$.
C.2 Prompts for Qwen-VL Evaluation
To further evaluate the performance of our method, we leverage Qwen-VL [62] for perceptual evaluation across three key aspects: age accuracy, identity consistency, and image quality. The prompts are structured as follows:
Age Estimation
"Please detect the age of the person in the image and return in the following format: age:{age}."
Image Quality
"Please evaluate image quality of the face image and provide a quality score (0–100), and return in the following format: quality:{quality_score}."
Identity Similarity
"Please evaluate if the individuals in these two images are the same person based solely on facial structure, ignoring factors such as style, lighting, age, or background. Provide a score between 1 (completely different) and 100 (completely identical), and return in the following format: similarity:{similarity_score}."
Each generated image is assessed using the corresponding prompt. When calculate identity similarity, the reference image is presented alongside the generated aged image to facilitate comparison. The resulting textual responses from Qwen-VL are parsed to extract quantitative scores.
C.3 Harmonic Consistency Score (HCS)
To jointly evaluate age accuracy and identity similarity, we introduce the Harmonic Consistency Score (HCS), defined as:
$A=\left(1-\frac{\text{MAE}}{M}\right)\cdot 100,\quad\text{HCS}=2\cdot\frac{A\cdot I}{A+I},$
where MAE denotes the mean absolute error between the predicted and target ages, and $M$ is the predefined maximum allowable age deviation (set to 40). The normalized age accuracy $A\in[0,100]$ reflects proximity to the target age, while $I\in[0,100]$ is the identity similarity score, obtained by multiplying the cosine similarity between ArcFace embeddings by 100. The harmonic formulation ensures a balanced evaluation that penalizes degradation in either attribute. Compared to simple averaging, the harmonic mean is more sensitive to low values, which is desirable in this context: a high HCS can only be achieved when both age accuracy and identity similarity are simultaneously high.
C.4 Open-Source Implementations and Settings
For reproducibility and comprehensive comparison, we evaluate several open-source face aging methods using their official pretrained models and inference pipelines. The following repositories are utilized:
Method Repository Link
SAM [1] https://github.com/yuval-alaluf/SAM
IPFE [3] https://github.com/sudban3089/ID-Preserving-Facial-Aging
FADING [6] https://github.com/MunchkinChen/FADING
CUSP [13] https://github.com/guillermogotre/CUSP
Lifespan [44] https://github.com/royorel/Lifespan_Age_Transformation_Synthesis
HRFAE [69] https://github.com/InterDigitalInc/HRFAE
Table 3: Open-source face aging methods and their official repositories.
All models are evaluated using standardized input settings and tested on the CelebA-HQ and CelebA-HQ (in-the-wild) test dataset. Due to the licensing terms of the CelebA [38] dataset , we are unable to display the original input images. Instead, we present the corresponding image inversions generated using null-text inversion [42]. We report metrics including age accuracy, identity similarity, image quality, and the proposed HCS. This unified evaluation protocol ensures fair and consistent performance comparison across diverse methods. For IPFE, which requires multiple images of the same identity as input, we randomly select one reference image per subject for identity similarity evaluation. Since both IPFE and FADING perform test-time tuning for each new input face image, a fixed training dataset is not applicable to these methods, and thus their training data size is not reported.
C.5 Age–ID Trade-off Evaluation Details
[…]
Figure 7: Age–ID trade-off curves of different methods. As the age shift value increases, either the Age cosine or ID cosine decreases for SAM, HRFAE, and IPFE. In contrast, our method maintains stable Age and ID consistency, showing no significant drop.
C.6 User Study Details
We conducted a user study to evaluate the perceptual quality of age-transformed face images generated by our method in comparison to two state-of-the-art methods, CUSP and SAM. A total of 50 volunteers participated in the survey. We randomly selected 20 identities from the CelebA-HQ test dataset and generated 6 age-progressed images for each, evenly spaced across ages ranging from 5 to 80 years. In each trial, participants were presented with pairs of image groups—each group consisting of the original reference face followed by the corresponding age-transformed images. The two groups (our method versus another method) were displayed side-by-side with randomized order to mitigate positional bias.
As illustrated in Fig. 12, the questionnaire provided clear guidance instructing participants to jointly assess three criteria: age realism, identity similarity, and overall visual quality. An example with labeled groups and comparative explanations was included to familiarize participants with the evaluation process. The study employed a forced-choice 1-vs-1 protocol, and results summarized in the main text demonstrate a consistent preference for our approach, confirming its superior ability to generate visually convincing and identity-preserving age transformations.
Appendix D Additional Results
D.1 Qualitative Comparison on in-the-wild Images
Fig. 8 presents a qualitative comparison of aging results on CelebA-HQ (in-the-wild) images, evaluating our method against state-of-the-art GAN-based and diffusion-based approaches. Compared to aligned datasets, in-the-wild images pose greater challenges due to diverse facial poses, complex backgrounds, occlusions, and uncontrolled lighting conditions. Under these challenging scenarios, the baseline methods exhibit various limitations. While CUSP and HRFAE maintain relatively high identity consistency, they often fail to capture realistic aging cues, resulting in over-smoothed faces with insufficient detail such as wrinkles and hair graying. LifeSpan, SAM, and Fading are prone to producing severe artifacts and significant identity drift, particularly under large age transformations, leading to unnatural facial structures and distorted textures. IPFE, on the other hand, generates faces with low visual fidelity and suffers from notable identity inconsistency, often producing blurry or distorted outputs. In contrast, our method demonstrates strong robustness and generalization in these real-world conditions, consistently generating high-fidelity aging results that preserve identity features while capturing fine-grained and realistic age-related changes such as wrinkle formation, hair graying, and structural facial transitions.
Figure 8: Qualitative comparison of aging results on CelebA-HQ (in-the-wild) images. Despite the challenges posed by real-world conditions such as occlusions, varying poses, and complex lighting and backgrounds, our method generates more photorealistic and coherent aging results, with better preservation of identity and more accurate aging effects including wrinkle formation, hair graying, and facial structure changes.
D.2 Generalization to Diverse Reference Ages
To thoroughly evaluate the age controllability and robustness of our method, we conduct face aging experiments using a wide range of reference ages, spanning from 1 to 80 years old. Specifically, we select reference images at 10-year intervals and generate aging results targeting six representative ages: 5, 20, 35, 50, 65, and 80, for each reference image. As shown in Fig. 11, our model demonstrates smooth and realistic age transformations across the entire age range, effectively handling both forward and backward aging transitions. The generated results exhibit consistent aging patterns, such as the gradual appearance of wrinkles, changes in skin texture, facial structural modifications, and hair graying, while preserving identity fidelity at each age target. These results highlight the strong generalization ability of our approach, ensuring effective age transformation across a variety of reference faces with diverse age inputs.
D.3 Face Aging across the Entire Lifespan
To further evaluate the age controllability of our approach, we conduct experiments generating human faces across the full age spectrum from 1 to 80 years. As shown in Fig. 9 and Fig. 10, our model produces smooth and continuous transitions of facial features across decades, accurately reflecting both age progression and regression. In contrast to prior works [3, 13, 69], which are typically limited to coarse age intervals (e.g., child, adult, elderly) or restricted age ranges, our method supports fine-grained age conditioning at each individual year without the need for additional retraining or manual tuning.
Figure 9: Face aging results from 1 to 80 years old. Reference images are marked in red.
Figure 10: Face aging results from 1 to 80 years old of in-the-wild images.
Figure 11: Aging results generated from diverse reference ages. For each reference image, we synthesize faces at six target ages: 5, 20, 35, 50, 65, and 80. Our method produces smooth and realistic age transitions across the entire lifespan, capturing both forward and backward aging effects while maintaining high identity consistency.
Figure 12: User study setup for comparing the visual outcomes of age-transformed face images. Participants were presented with pairs of image groups, each showing the original face alongside age-transformed images from our method and the baselines (CUSP and SAM). They assessed three criteria: age realism, identity similarity, and overall visual quality.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See Abstract and Introduction (section 1).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Limitations (Appendix A).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: See Method (section 3).
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See Experiments (section 4) and Experiments details (Appendix C).
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: Due to the privacy concerns associated with facial data and company policy restrictions, the full dataset and complete training code cannot be released at this time. However, we plan to release the dataset and code in the future to support community use and facilitate reproducibility.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See Experiments (section 4) and Experiments details (Appendix C).
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: See Experiments (section 4).
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Experiments (section 4) and Experiments details (Appendix C).
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have carefully checked the NeurIPS code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: see Broader Impacts (Appendix B).
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: see Broader Impacts (Appendix B).
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We politely cited the existing assets and mentioned the license and terms of use in Experiments (section 4) and Experiments details (Appendix C).
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper dose not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: See User Study Details (Appendix C.6).
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 This significantly broadens its applicability to more diverse and unconstrained real-world scenarios.
+Code is available at https://github.com/byliutao/Cradle2Cane.
 1 Introduction
@@
 To this end, numerous methods have been developed to achieve high-quality, identity-preserving age progression and regression.
-Recent approaches are based on deep generative models, such as generative adversarial networks (GANs) alaluf2021only ; goodfellow2014gan ; hsu2022agetransgan and Diffusion Models (DMs) banerjee2023identity ; chen2023face ; ito2025selfage ; wahid2024diffage3d , and have shown promising results.
+Recent approaches are based on deep generative models, such as generative adversarial networks (GANs) alaluf2021only ; goodfellow2014gan ; hsu2022agetransgan and Diffusion Models (DMs) banerjee2023identity ; chen2023face ; ito2025selfage ; wahid2024diffage3d ; wu2025representationentanglementgenerationtraining , and have shown promising results.
 However, to the best of our knowledge, most existing methods suffer from a limited transformation range and often struggle to maintain high-fidelity results when handling large age gaps, occlusions, or extreme head poses.
@@
 In contrast, lower noise levels better preserve identity information with less aging accuracy, highlighting the trade-off between visual age change and identity consistency.
-Similar identity-editing trade-offs are also observed in image editing and generation hertz2022prompt ; li2023stylediffusion ; meng2021sdedit ; kai2023DPL .
-However, directly applying the few-step diffusion models cannot achieve fine-grained face aging with identity consistency and age accuracy, resulting from that the few-step T2I diffusion models ho2022imagen ; Rombach_2022_CVPR_stablediffusion ; deepfloyd do not inherently support age or identity conditions.
+Similar identity-editing trade-offs are also observed in image editing and generation hertz2022prompt ; li2023stylediffusion ; meng2021sdedit ; kai2023DPL ; tang2024locinv ; hu2025anchor_token .
+However, directly applying the few-step diffusion models cannot achieve fine-grained face aging with identity consistency and age accuracy, resulting from that the few-step T2I diffusion models ho2022imagen ; Rombach_2022_CVPR_stablediffusion ; deepfloyd ; li2025oneway_ticket do not inherently support age or identity conditions.
 Figure 1: Age–ID trade-off curves across sixty age shift values.
@@
 Best results are marked in blue, and second-best in green.
-Method Type Face++ Evaluation (CelebA-HQ) Qwen-VL Evaluation (CelebA-HQ) Qwen-VL (CelebA-HQ-in-the-wild) Inference Time (s) Train Data
+Method Type Face++ Evaluation (CelebA-HQ) Qwen-VL Evaluation (CelebA-HQ) Qwen-VL (CelebA-HQ-in-the-wild)
+Inference
+Time (s)
+Train
+Data
 Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑ Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑ Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑
@@
 Figure 6: User Study
-Ours vs CUSPOurs vs SAM0025252525505050507575757510010010010072727272676767672828282833333333
+Ours vs CUSPOurs vs SAM02525505075751001007272676728283333
 User Study.
@@
 In addition, Cradle2Cane demonstrates strong generalization to real-world scenarios by effectively handling in-the-wild human face images, a setting where existing methods often fail.
-Appendix
-Appendix A Limitations
-While our method achieves state-of-the-art performance in balancing age realism and identity consistency, there remain several limitations that merit discussion.
-In cases of extreme age transformation (e.g., from a child to an elderly person or vice versa), the model tends to favor facial realism and age accuracy at the cost of preserving some visual details in the original image.
-For example, accessories such as eyeglasses, earrings, or clothing color may not always be faithfully retained after editing, as they are not explicitly modeled or enforced during training.
-This issue stems from the adaptive noise injection design, which purposefully increases editability for large age gaps, potentially altering finer image semantics beyond facial identity.
-Appendix B Broader Impacts
-Our proposed face aging method provides a flexible framework for independently controlling visual age via a two-pass diffusion process.
-This enables a range of positive applications, including digital entertainment (e.g., age effects in movies or games), age-invariant face recognition, and future appearance simulation for healthcare or counseling.
-All experiments are conducted on anonymized public benchmarks under ethical research settings.
-At the same time, we recognize the potential risks associated with misuse.
-The ability to generate photorealistic age manipulation with identity consistency may facilitate malicious uses such as identity spoofing, misinformation, or privacy violations.
-We strongly discourage unauthorized or commercial deployment without safeguards like watermarking, traceable provenance, or human review.
-We hope this work inspires further progress toward ethical and responsible generative modeling in the vision community.
-Appendix C Experiments details
-C.1 Implementation Details.
-[…]
-The U-Net backbone of SDXL-Turbo is fine-tuned via LoRA, and training is conducted on 8 NVIDIA A6000 GPUs with a batch size of 4.
-For hyperparameters, in SVR-ArcFace, we set $\alpha=0.01$ and $\beta=1.2$, while in Rotate-CLIP, we use an interpolation strength of $\lambda=0.5$.
-For the training loss terms, we set the weights as $\lambda_{1}=0.25$, $\lambda_{2}=1.2$, $\lambda_{3}=1.5$, $\lambda_{4}=1.5$, $\lambda_{5}=0.25$, and $\lambda_{6}=0.1$.
-C.2 Prompts for Qwen-VL Evaluation
-To further evaluate the performance of our method, we leverage Qwen-VL [62] for perceptual evaluation across three key aspects: age accuracy, identity consistency, and image quality.
-The prompts are structured as follows:
-Age Estimation
-"Please detect the age of the person in the image and return in the following format: age:{age}."
-Image Quality
-"Please evaluate image quality of the face image and provide a quality score (0–100), and return in the following format: quality:{quality_score}."
-Identity Similarity
-"Please evaluate if the individuals in these two images are the same person based solely on facial structure, ignoring factors such as style, lighting, age, or background.
-Provide a score between 1 (completely different) and 100 (completely identical), and return in the following format: similarity:{similarity_score}."
-Each generated image is assessed using the corresponding prompt.
-When calculate identity similarity, the reference image is presented alongside the generated aged image to facilitate comparison.
-The resulting textual responses from Qwen-VL are parsed to extract quantitative scores.
-C.3 Harmonic Consistency Score (HCS)
-To jointly evaluate age accuracy and identity similarity, we introduce the Harmonic Consistency Score (HCS), defined as:
-$A=\left(1-\frac{\text{MAE}}{M}\right)\cdot 100,\quad\text{HCS}=2\cdot\frac{A\cdot I}{A+I},$
-where MAE denotes the mean absolute error between the predicted and target ages, and $M$ is the predefined maximum allowable age deviation (set to 40).
-The normalized age accuracy $A\in[0,100]$ reflects proximity to the target age, while $I\in[0,100]$ is the identity similarity score, obtained by multiplying the cosine similarity between ArcFace embeddings by 100.
-The harmonic formulation ensures a balanced evaluation that penalizes degradation in either attribute.
-Compared to simple averaging, the harmonic mean is more sensitive to low values, which is desirable in this context: a high HCS can only be achieved when both age accuracy and identity similarity are simultaneously high.
-C.4 Open-Source Implementations and Settings
-For reproducibility and comprehensive comparison, we evaluate several open-source face aging methods using their official pretrained models and inference pipelines.
-The following repositories are utilized:
-Method Repository Link
-SAM [1] https://github.com/yuval-alaluf/SAM
-IPFE [3] https://github.com/sudban3089/ID-Preserving-Facial-Aging
-FADING [6] https://github.com/MunchkinChen/FADING
-CUSP [13] https://github.com/guillermogotre/CUSP
-Lifespan [44] https://github.com/royorel/Lifespan_Age_Transformation_Synthesis
-HRFAE [69] https://github.com/InterDigitalInc/HRFAE
-Table 3: Open-source face aging methods and their official repositories.
-All models are evaluated using standardized input settings and tested on the CelebA-HQ and CelebA-HQ (in-the-wild) test dataset.
-Due to the licensing terms of the CelebA [38] dataset , we are unable to display the original input images.
-Instead, we present the corresponding image inversions generated using null-text inversion [42].
-We report metrics including age accuracy, identity similarity, image quality, and the proposed HCS.
-This unified evaluation protocol ensures fair and consistent performance comparison across diverse methods.
-For IPFE, which requires multiple images of the same identity as input, we randomly select one reference image per subject for identity similarity evaluation.
-Since both IPFE and FADING perform test-time tuning for each new input face image, a fixed training dataset is not applicable to these methods, and thus their training data size is not reported.
-C.5 Age–ID Trade-off Evaluation Details
-[…]
-Figure 7: Age–ID trade-off curves of different methods.
-As the age shift value increases, either the Age cosine or ID cosine decreases for SAM, HRFAE, and IPFE.
-In contrast, our method maintains stable Age and ID consistency, showing no significant drop.
-C.6 User Study Details
-We conducted a user study to evaluate the perceptual quality of age-transformed face images generated by our method in comparison to two state-of-the-art methods, CUSP and SAM.
-A total of 50 volunteers participated in the survey.
-We randomly selected 20 identities from the CelebA-HQ test dataset and generated 6 age-progressed images for each, evenly spaced across ages ranging from 5 to 80 years.
-In each trial, participants were presented with pairs of image groups—each group consisting of the original reference face followed by the corresponding age-transformed images.
-The two groups (our method versus another method) were displayed side-by-side with randomized order to mitigate positional bias.
-As illustrated in Fig.
-12, the questionnaire provided clear guidance instructing participants to jointly assess three criteria: age realism, identity similarity, and overall visual quality.
-An example with labeled groups and comparative explanations was included to familiarize participants with the evaluation process.
-The study employed a forced-choice 1-vs-1 protocol, and results summarized in the main text demonstrate a consistent preference for our approach, confirming its superior ability to generate visually convincing and identity-preserving age transformations.
-Appendix D Additional Results
-D.1 Qualitative Comparison on in-the-wild Images
-Fig.
-8 presents a qualitative comparison of aging results on CelebA-HQ (in-the-wild) images, evaluating our method against state-of-the-art GAN-based and diffusion-based approaches.
-Compared to aligned datasets, in-the-wild images pose greater challenges due to diverse facial poses, complex backgrounds, occlusions, and uncontrolled lighting conditions.
-Under these challenging scenarios, the baseline methods exhibit various limitations.
-While CUSP and HRFAE maintain relatively high identity consistency, they often fail to capture realistic aging cues, resulting in over-smoothed faces with insufficient detail such as wrinkles and hair graying.
-LifeSpan, SAM, and Fading are prone to producing severe artifacts and significant identity drift, particularly under large age transformations, leading to unnatural facial structures and distorted textures.
-IPFE, on the other hand, generates faces with low visual fidelity and suffers from notable identity inconsistency, often producing blurry or distorted outputs.
-In contrast, our method demonstrates strong robustness and generalization in these real-world conditions, consistently generating high-fidelity aging results that preserve identity features while capturing fine-grained and realistic age-related changes such as wrinkle formation, hair graying, and structural facial transitions.
-Figure 8: Qualitative comparison of aging results on CelebA-HQ (in-the-wild) images.
-Despite the challenges posed by real-world conditions such as occlusions, varying poses, and complex lighting and backgrounds, our method generates more photorealistic and coherent aging results, with better preservation of identity and more accurate aging effects including wrinkle formation, hair graying, and facial structure changes.
-D.2 Generalization to Diverse Reference Ages
-To thoroughly evaluate the age controllability and robustness of our method, we conduct face aging experiments using a wide range of reference ages, spanning from 1 to 80 years old.
-Specifically, we select reference images at 10-year intervals and generate aging results targeting six representative ages: 5, 20, 35, 50, 65, and 80, for each reference image.
-As shown in Fig.
-11, our model demonstrates smooth and realistic age transformations across the entire age range, effectively handling both forward and backward aging transitions.
-The generated results exhibit consistent aging patterns, such as the gradual appearance of wrinkles, changes in skin texture, facial structural modifications, and hair graying, while preserving identity fidelity at each age target.
-These results highlight the strong generalization ability of our approach, ensuring effective age transformation across a variety of reference faces with diverse age inputs.
-D.3 Face Aging across the Entire Lifespan
-To further evaluate the age controllability of our approach, we conduct experiments generating human faces across the full age spectrum from 1 to 80 years.
-As shown in Fig.
-9 and Fig.
-10, our model produces smooth and continuous transitions of facial features across decades, accurately reflecting both age progression and regression.
-In contrast to prior works [3, 13, 69], which are typically limited to coarse age intervals (e.g., child, adult, elderly) or restricted age ranges, our method supports fine-grained age conditioning at each individual year without the need for additional retraining or manual tuning.
-Figure 9: Face aging results from 1 to 80 years old. Reference images are marked in red.
-Figure 10: Face aging results from 1 to 80 years old of in-the-wild images.
-Figure 11: Aging results generated from diverse reference ages.
-For each reference image, we synthesize faces at six target ages: 5, 20, 35, 50, 65, and 80.
-Our method produces smooth and realistic age transitions across the entire lifespan, capturing both forward and backward aging effects while maintaining high identity consistency.
-Figure 12: User study setup for comparing the visual outcomes of age-transformed face images.
-Participants were presented with pairs of image groups, each showing the original face alongside age-transformed images from our method and the baselines (CUSP and SAM).
-They assessed three criteria: age realism, identity similarity, and overall visual quality.
+Acknowledgements
+This research was supported by the collaborative project between Beijing Samsung Telecommunications Technology Co., Ltd.
+and the Tianjin Key Laboratory of Visual Computing and Intelligent Perception (VCIP) of Nankai University, entitled “Generated Face for Enhancing Face Dataset” (Project No.
+SRC-Beijing-DVL-2024-00241).
+We would like to express our sincere gratitude to all co-authors for their invaluable contributions and insightful suggestions.
+We are particularly grateful to Yaxing Wang (Associate Professor, Nankai University) and Kai Wang (Assistant Professor, City University of Hong Kong (Dongguan)).
+Their meticulous advice and guidance were instrumental in the completion of this research.
 NeurIPS Paper Checklist
@@
     • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix
+Appendix A Limitations
+While our method achieves state-of-the-art performance in balancing age realism and identity consistency, there remain several limitations that merit discussion.
+In cases of extreme age transformation (e.g., from a child to an elderly person or vice versa), the model tends to favor facial realism and age accuracy at the cost of preserving some visual details in the original image.
+For example, accessories such as eyeglasses, earrings, or clothing color may not always be faithfully retained after editing, as they are not explicitly modeled or enforced during training.
+This issue stems from the adaptive noise injection design, which purposefully increases editability for large age gaps, potentially altering finer image semantics beyond facial identity.
+Appendix B Broader Impacts
+Our proposed face aging method provides a flexible framework for independently controlling visual age via a two-pass diffusion process.
+This enables a range of positive applications, including digital entertainment (e.g., age effects in movies or games), age-invariant face recognition, and future appearance simulation for healthcare or counseling.
+All experiments are conducted on anonymized public benchmarks under ethical research settings.
+At the same time, we recognize the potential risks associated with misuse.
+The ability to generate photorealistic age manipulation with identity consistency may facilitate malicious uses such as identity spoofing, misinformation, or privacy violations.
+We strongly discourage unauthorized or commercial deployment without safeguards like watermarking, traceable provenance, or human review.
+We hope this work inspires further progress toward ethical and responsible generative modeling in the vision community.
+Appendix C Experiments details
+C.1 Implementation Details.
+Our framework is built upon the SDXL-Turbo architecture and trained on the FFHQ dataset, which contains high-quality facial images annotated with age and gender labels.
+To improve background consistency, we employ CarveKit [60] for foreground-background segmentation, replacing non-facial regions with a uniform gray mask.
+During training, we guide the model using text prompts in the format: "a face image of a {target age} years old {gender}" where gender is either female or male.
+All images are generated at a fixed resolution of $\times512$.
+The U-Net backbone of SDXL-Turbo is fine-tuned via LoRA, and training is conducted on 8 NVIDIA A6000 GPUs with a batch size of 4.
+For hyperparameters, in SVR-ArcFace, we set $\alpha=0.01$ and $\beta=1.2$, while in Rotate-CLIP, we use an interpolation strength of $\lambda=0.5$.
+For the training loss terms, we set the weights as $\lambda_{1}=0.25$, $\lambda_{2}=1.2$, $\lambda_{3}=1.5$, $\lambda_{4}=1.5$, $\lambda_{5}=0.25$, and $\lambda_{6}=0.1$.
+C.2 Prompts for Qwen-VL Evaluation
+To further evaluate the performance of our method, we leverage Qwen-VL [68] for perceptual evaluation across three key aspects: age accuracy, identity consistency, and image quality.
+The prompts are structured as follows:
+Age Estimation
+"Please detect the age of the person in the image and return in the following format: age:{age}."
+Image Quality
+"Please evaluate image quality of the face image and provide a quality score (0–100), and return in the following format: quality:{quality_score}."
+Identity Similarity
+"Please evaluate if the individuals in these two images are the same person based solely on facial structure, ignoring factors such as style, lighting, age, or background.
+Provide a score between 1 (completely different) and 100 (completely identical), and return in the following format: similarity:{similarity_score}."
+Each generated image is assessed using the corresponding prompt.
+When calculate identity similarity, the reference image is presented alongside the generated aged image to facilitate comparison.
+The resulting textual responses from Qwen-VL are parsed to extract quantitative scores.
+C.3 Harmonic Consistency Score (HCS)
+To jointly evaluate age accuracy and identity similarity, we introduce the Harmonic Consistency Score (HCS), defined as:
+$A=\left(1-\frac{\text{MAE}}{M}\right)\cdot 100,\quad\text{HCS}=2\cdot\frac{A\cdot I}{A+I},$
+where MAE denotes the mean absolute error between the predicted and target ages, and $M$ is the predefined maximum allowable age deviation (set to 40).
+The normalized age accuracy $A\in[0,100]$ reflects proximity to the target age, while $I\in[0,100]$ is the identity similarity score, obtained by multiplying the cosine similarity between ArcFace embeddings by 100.
+The harmonic formulation ensures a balanced evaluation that penalizes degradation in either attribute.
+Compared to simple averaging, the harmonic mean is more sensitive to low values, which is desirable in this context: a high HCS can only be achieved when both age accuracy and identity similarity are simultaneously high.
+C.4 Open-Source Implementations and Settings
+For reproducibility and comprehensive comparison, we evaluate several open-source face aging methods using their official pretrained models and inference pipelines.
+The following repositories are utilized:
+Method Repository Link
+SAM [1] https://github.com/yuval-alaluf/SAM
+IPFE [3] https://github.com/sudban3089/ID-Preserving-Facial-Aging
+FADING [7] https://github.com/MunchkinChen/FADING
+CUSP [14] https://github.com/guillermogotre/CUSP
+Lifespan [48] https://github.com/royorel/Lifespan_Age_Transformation_Synthesis
+HRFAE [76] https://github.com/InterDigitalInc/HRFAE
+Table 3: Open-source face aging methods and their official repositories.
+All models are evaluated using standardized input settings and tested on the CelebA-HQ and CelebA-HQ (in-the-wild) test dataset.
+Due to the licensing terms of the CelebA [41] dataset , we are unable to display the original input images.
+Instead, we present the corresponding image inversions generated using null-text inversion [45].
+We report metrics including age accuracy, identity similarity, image quality, and the proposed HCS.
+This unified evaluation protocol ensures fair and consistent performance comparison across diverse methods.
+For IPFE, which requires multiple images of the same identity as input, we randomly select one reference image per subject for identity similarity evaluation.
+Since both IPFE and FADING perform test-time tuning for each new input face image, a fixed training dataset is not applicable to these methods, and thus their training data size is not reported.
+C.5 Age–ID Trade-off Evaluation Details
+To quantitatively evaluate the trade-off between age accuracy and identity consistency, we selected 50 male and 50 female face images from the CelebA-HQ test dataset.
+For each image, we generated aging results across age offsets ranging from 60-60 to +60+60 with a step size of $1$ year, excluding the zero offset.
+Identity similarity was measured using ArcFace by computing the cosine similarity between the original and age-edited images.
+Age similarity was quantified based on the predicted age error from the MiVOLO estimator.
+Specifically, we computed the age consistency score as $\text{age\_cosine}=1-\frac{|\text{age}_{\text{pred}}-\text{age}_{\text{target}}|}{\text{max\_age\_diff}}$, where max_age_diff is set to $40$.
+This score ranges from 0 to $1$, with higher values indicating better age alignment.
+Comparisons with the remaining methods are illustrated in Fig.
+7.
+Note that IPFE [3] does not support continuous year-level age control, thus only the provided age groups reported in the original paper were included in our evaluation.
+Figure 7: Age–ID trade-off curves of different methods.
+As the age shift value increases, either the Age cosine or ID cosine decreases for SAM, HRFAE, and IPFE.
+In contrast, our method maintains stable Age and ID consistency, showing no significant drop.
+C.6 User Study Details
+We conducted a user study to evaluate the perceptual quality of age-transformed face images generated by our method in comparison to two state-of-the-art methods, CUSP and SAM.
+A total of 50 volunteers participated in the survey.
+We randomly selected 20 identities from the CelebA-HQ test dataset and generated 6 age-progressed images for each, evenly spaced across ages ranging from 5 to 80 years.
+In each trial, participants were presented with pairs of image groups—each group consisting of the original reference face followed by the corresponding age-transformed images.
+The two groups (our method versus another method) were displayed side-by-side with randomized order to mitigate positional bias.
+As illustrated in Fig.
+12, the questionnaire provided clear guidance instructing participants to jointly assess three criteria: age realism, identity similarity, and overall visual quality.
+An example with labeled groups and comparative explanations was included to familiarize participants with the evaluation process.
+The study employed a forced-choice 1-vs-1 protocol, and results summarized in the main text demonstrate a consistent preference for our approach, confirming its superior ability to generate visually convincing and identity-preserving age transformations.
+Appendix D Additional Results
+D.1 Qualitative Comparison on in-the-wild Images
+Fig.
+8 presents a qualitative comparison of aging results on CelebA-HQ (in-the-wild) images, evaluating our method against state-of-the-art GAN-based and diffusion-based approaches.
+Compared to aligned datasets, in-the-wild images pose greater challenges due to diverse facial poses, complex backgrounds, occlusions, and uncontrolled lighting conditions.
+Under these challenging scenarios, the baseline methods exhibit various limitations.
+While CUSP and HRFAE maintain relatively high identity consistency, they often fail to capture realistic aging cues, resulting in over-smoothed faces with insufficient detail such as wrinkles and hair graying.
+LifeSpan, SAM, and Fading are prone to producing severe artifacts and significant identity drift, particularly under large age transformations, leading to unnatural facial structures and distorted textures.
+IPFE, on the other hand, generates faces with low visual fidelity and suffers from notable identity inconsistency, often producing blurry or distorted outputs.
+In contrast, our method demonstrates strong robustness and generalization in these real-world conditions, consistently generating high-fidelity aging results that preserve identity features while capturing fine-grained and realistic age-related changes such as wrinkle formation, hair graying, and structural facial transitions.
+Figure 8: Qualitative comparison of aging results on CelebA-HQ (in-the-wild) images.
+Despite the challenges posed by real-world conditions such as occlusions, varying poses, and complex lighting and backgrounds, our method generates more photorealistic and coherent aging results, with better preservation of identity and more accurate aging effects including wrinkle formation, hair graying, and facial structure changes.
+D.2 Generalization to Diverse Reference Ages
+To thoroughly evaluate the age controllability and robustness of our method, we conduct face aging experiments using a wide range of reference ages, spanning from 1 to 80 years old.
+Specifically, we select reference images at 10-year intervals and generate aging results targeting six representative ages: 5, 20, 35, 50, 65, and 80, for each reference image.
+As shown in Fig.
+11, our model demonstrates smooth and realistic age transformations across the entire age range, effectively handling both forward and backward aging transitions.
+The generated results exhibit consistent aging patterns, such as the gradual appearance of wrinkles, changes in skin texture, facial structural modifications, and hair graying, while preserving identity fidelity at each age target.
+These results highlight the strong generalization ability of our approach, ensuring effective age transformation across a variety of reference faces with diverse age inputs.
+D.3 Face Aging across the Entire Lifespan
+To further evaluate the age controllability of our approach, we conduct experiments generating human faces across the full age spectrum from 1 to 80 years.
+As shown in Fig.
+9 and Fig.
+10, our model produces smooth and continuous transitions of facial features across decades, accurately reflecting both age progression and regression.
+In contrast to prior works [3, 14, 76], which are typically limited to coarse age intervals (e.g., child, adult, elderly) or restricted age ranges, our method supports fine-grained age conditioning at each individual year without the need for additional retraining or manual tuning.
+Appendix E Additional Experiment
+E.1 Additional Quantitative Experiment
+To further validate the effectiveness of our approach, we conducted additional quantitative experiments on extra datasets and baselines.
+Specifically, we compared our method with recent generic face-editing systems (e.g., StyleCLIP [50] and FaceDNeRF [78]) on the standard aging datasets AgeDB-30 [46], CACD [6], and FG-NET.
+For each dataset, we generated approximately 200 images per method and evaluated them using four metrics: age difference (Age Diff.), identity similarity (ID Sim.), image quality (Img.
+Quality), and Holistic Consistency Scores (HCS).
+The comprehensive results are summarized in Table 4, where lower Age Diff.
+and higher scores on the remaining metrics indicate better performance.
+Method AgeDB-30 CACD FG-NET
+Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑ Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑ Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑
+CUSP [14] 8.29 76.30 82.80 77.75 10.80 75.40 85.86 74.18 19.45 74.20 79.79 60.71
+SAM [1] 8.72 58.91 86.57 67.19 9.25 67.27 87.57 71.75 5.33 72.18 81.98 78.77
+FADING [7] 8.10 76.00 79.22 77.82 5.16 71.27 86.44 78.39 7.22 73.15 79.43 77.30
+Styleclip [50] 14.73 62.12 86.60 62.64 16.28 70.64 87.80 64.48 14.73 73.82 88.19 68.08
+Facednerf [78] 12.60 53.78 91.97 60.25 13.05 59.00 92.02 62.91 13.21 56.37 92.43 61.21
+Cradle2Cane 5.57 73.00 90.01 79.00 4.63 66.73 90.02 76.06 5.79 67.18 85.03 75.25
+Table 4: Comparison on AgeDB-30, CACD and FG-NET datasets. Best results are marked in blue, and second-best in green.
+Our method, Cradle2Cane, demonstrates exceptional performance across all three benchmarks.
+On AgeDB-30, it achieves state-of-the-art results, securing the best age accuracy and the top harmonic score, which underscores its superior overall balance.
+This strong performance continues on the CACD dataset, where our method again delivers the best age accuracy while maintaining highly competitive scores on other metrics.
+Even on the highly challenging FG-NET dataset, Cradle2Cane proves its robustness by delivering the second-best age accuracy and demonstrating strong, consistent performance against all competitors.
+E.2 Ablation Study on Threshold Selection
+To investigate the rationale behind our choice of age thresholds, we conducted an ablation study with multiple division settings.
+Specifically, we compared thresholds of {5,15}, {5,20}, {7,22}, {10,30}, {15,35}, and {20,40}.
+The results presented in Table 5 reveal a distinct trade-off between age accuracy and identity preservation.
+We observe that wider threshold intervals, such as {15,35} and {20,40}, yield superior identity similarity (ID Sim.) and Holistic Consistency Scores (HCS).
+However, this gain is achieved at the expense of age fidelity, as evidenced by the significant increase in the Age Difference metric from a low of 4.92 to 6.71.
+Given that age accuracy is the central objective of our work, we selected the {5, 20}division as our default configuration.
+This setting achieves the best performance in age accuracy while maintaining a competitive balance in identity preservation and image quality.
+This choice ensures our primary goal is met and establishes a rigorous, transparent baseline for our experiments.
+Table 5: Ablation study on different threshold divisions.
+This analysis highlights the trade-off between age accuracy and identity preservation.
+Best results are marked in blue, and second-best in green.
+Thresholds Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑
+{5, 15} 4.93 73.18 92.28 79.77
+{5, 20} 4.92 73.73 92.60 80.11
+{7, 22} 5.06 74.82 92.89 80.60
+{10, 30} 5.46 76.91 92.56 81.36
+{15, 35} 5.64 77.09 92.82 81.26
+{20, 40} 6.71 78.55 92.74 80.82
+E.3 Ablation Study on Robustness and Architectural Contributions
+To verify whether the robustness of our method on in-the-wild face images primarily comes from background removal pre-processing (Carvekit) or from the model architecture itself, we conducted an ablation study.
+Specifically, we compared our method with and without Carvekit pre-processing.
+Both variants were trained for 10 epochs on 1,000 images and evaluated on the CelebA-in-the-wild dataset.
+The results are shown in Table 6.
+Table 6: Ablation study of the background removal pre-processing (Carvekit).
+The minor performance difference highlights that robustness is intrinsic to the model architecture.
+Best results are marked in blue.
+Method Age Diff. ↓ ID Sim. ↑ Img. Quality ↑ HCS ↑
+w/ Carvekit 6.76 62.00 91.75 71.02
+w/o Carvekit 7.00 62.27 91.60 70.97
+As presented in the table, the performance impact of background removal is marginal.
+This finding indicates that Carvekit serves as a beneficial but non-essential preprocessing step, rather than the primary source of the model’s robustness.
+Instead, the method’s resilience to in-the-wild variations is primarily attributed to its architectural design.
+First, the SDXL-Turbo backbone, pre-trained on a vast and diverse dataset, provides a strong foundation for generalization across varied poses, lighting conditions, and expressions.
+Second, our proposed IDEmb module systematically reinforces identity preservation.
+The SVR-ArcFace component extracts a stable identity embedding that is disentangled from transient attributes, while Rotate-CLIP executes minimal, precise modifications within CLIP’s robust semantic space.
+This dual mechanism ensures that the age attribute is altered while other original characteristics are preserved with high fidelity.
+In summary, this study confirms that our model’s robustness is an intrinsic property of its architecture, with SDXL-Turbo enabling generalization and IDEmb ensuring identity-consistent editing.
+Algorithm 1 The Proposed Cradle2Cane Framework
+1:Input: Source image $\mathbf{x}_{a}$, source age $a$, target age $b$.
+2:Output: Final aged image $\mathbf{x}_{b}$.
+3:
+4:// Pass 1: Adaptive Age Transformation
+5:Select noise level $\mathbf{z}_{i}$ based on $|\Delta\text{age}|$ per (Eq. 1).
+[…]
+7:Denoise from noise level $\mathbf{z}_{i}$ with condition $c_{\text{age}}$ to get latent $\hat{\mathbf{z}}_{0}$.
+[…]
+9:
+10:// Pass 2: Identity Enhancement
+11:// — IDEmb Generation —
+12:Generate aged variants $\{\mathbf{x}_{b}^{(i)}\}_{i=1}^{n}$ from $\mathbf{x}_{a}$.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+19:
+[…]
+[…]
+[…]
+23:
+24:// — Final Refinement —
+[…]
+[…]
+27:Set low noise timestep $T_{\text{low}}$.
+[…]
+29:for $t=T_{\text{low}},\dots,1$ do $\mathbf{z}_{t-1}\leftarrow p_{\theta}(\mathbf{z}_{t},t,c_{\text{ID}})$ $\triangleright$ Denoise with identity condition
+30:end for
+[…]
+32:
+33:Return $\mathbf{x}_{b}$.
+Figure 9: Face aging results from 1 to 80 years old. Reference images are marked in red.
+Figure 10: Face aging results from 1 to 80 years old of in-the-wild images.
+Figure 11: Aging results generated from diverse reference ages.
+For each reference image, we synthesize faces at six target ages: 5, 20, 35, 50, 65, and 80.
+Our method produces smooth and realistic age transitions across the entire lifespan, capturing both forward and backward aging effects while maintaining high identity consistency.
+Figure 12: User study setup for comparing the visual outcomes of age-transformed face images.
+Participants were presented with pairs of image groups, each showing the original face alongside age-transformed images from our method and the baselines (CUSP and SAM).
+They assessed three criteria: age realism, identity similarity, and overall visual quality.